# Mitigating Tail Narrowing in LLM Self-Improvement via Socratic-Guided Sampling

### v1

@@ Abstract
Yiwen Ding1thanks:  Equal contribution.thanks:  Work done during internship at Meituan., Zhiheng Xi [^1] footnotemark: 1, Wei He1, Zhuoyuan Li3, Yitao Zhai2, Xiaowei Shi2, Xunliang Cai2, Tao Gui1thanks:  Corresponding authors., Qi Zhang1, Xuanjing Huang133footnotemark: 3 1 Fudan University   2 Meituan 3 Macau University of Science and Technology {ywding23,zhxi22}@m.fudan.edu.cn, {tgui, xjhuang}@fudan.edu.cn
@@ 1 Introduction
To address this imbalance, a common approach is to allocate more sampling budget to the under-sampled, challenging queries (Tong et al., 2024), but this can be much more costly. In this paper, we propose an efficient and effective method called Guided Self-Improvement (GSI), which employs Socratic-style guidance signals (Chang, 2023; Dong et al., 2023b) to assist language models in exploring solutions for challenging queries. Specifically, we introduce an extra resampling phase called distribution re-balancing for difficult queries, applied after the generation step in the self-improvement process. During this phase, we provide targeted guidance signals to reduce sampling difficulty, narrow the sampling space, and minimize hallucinations during reasoning (Xie et al., 2023; Xi et al., 2024). As a result, GSI improves sampling quality, increases solution coverage (Bansal et al., 2024) for challenging queries, and mitigates the issue of tail narrowing.
@@ 2.1 Self-improvement for LLMs
Once filtered, the high-quality data supports post-training through methods like SFT (Zelikman et al., 2022) or preference-based techniques like Direct Preference Optimization (DPO, Yuan et al., 2024). This process is often iterative, allowing models to continually generate new data, filter it, and use it to refine their performance further (Zelikman et al., 2022; Gülçehre et al., 2023; Yuan et al., 2024).
@@ 3.1 Formulation of Self-improvement
Given a large language model $M_{0}$ and the original training dataset $\mathcal{D}=\{(x_{i},r_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is the problem, $r_{i}$ is the chain-of-thought rationale (Wei et al., 2022b) and $y_{i}$ represents the final answer. Each rationale $r_{i}$ consists of several intermediate steps, i.e., $r_{i}=[r_{i,1},\dots,r_{i,L}]$. The self-improvement process iteratively leverages the model’s correct responses to enhance its problem-solving abilities gradually. The process involves $T$ iterations, where each iteration comprises two main steps: Generate and Improve.
@@
At iteration $t$, based on the previous model $M_{t-1}$, we perform $K$ sampling operations for each query $x_{i}\in\mathcal{D}$:
@@
The newly generated data points form a set $\mathcal{D}^{{}^{\prime}}=\{(x_{i},\hat{r}_{i}^{j},\hat{y}_{i}^{j})\mid x_{i}\in\mathcal{D},j=[1,K]\}$. Each sampled $\hat{y_{i}}$ is evaluated using a reward function $\textrm{rf}(y_{i},\hat{y_{i}})\in\{0,1\}$, which checks its correctness based on ground-truth answers $y_{i}$. Only the correct solutions are filtered to form the high-quality dataset $\mathcal{D}_{t}$ at iteration $t$.
@@
In the $t$-th iteration, we use the self-generated high-quality dataset $\mathcal{D}_{t}$ to fine-tune the model $M_{t}$. The fine-tuning objective is to minimize the negative log-likelihood (NLL) loss:
@@ 4 Performance Bottleneck and Tail Narrowing in Self-Improvement
Despite extensive research into self-improvement methods (Gülçehre et al., 2023; Singh et al., 2024), the performance dynamics across successive training iterations remain underexplored. To this end, we conduct experiments across four backbone models to investigate the effects of sampling and tail data during iterative training.
@@
To uncover the relationship between performance trends with sampling, we vary the number of sampling $k$. Our results indicate that lower sampling times (e.g., $k=1$ or $k=2$) degrade model performance, leading to negative gains. Notably, as depicted by the purple solid lines in Figure 2, Llama2-7B, Mistral-7B, and Llama3-8B underperform compared to SFT on the original dataset (corresponding to the performance of iteration 1). This decline stems from the weaker reasoning abilities of these models, which limit the coverage of challenging queries. Consequently, the models gradually tackle only basic problems, resulting in a degradation or collapse of reasoning ability (Shumailov et al., 2023; Dohmatob et al., 2024a). As $k$ increases, the most notable performance gains occur during the first iteration, shown by the green and orange solid lines in Figure 2. However, after the third iteration, the progress halts, eventually reaching a performance plateau.
@@
As previously discussed, merely scaling the number of sampling eventually encounters performance bottlenecks. However, the low-probability and challenging tail data are often considered more crucial for improving model performance (Sorscher et al., 2022; Liu et al., 2024; Tong et al., 2024). To investigate the impact of tail data, we conduct additional experiments targeting these difficult examples. For queries that do not yield a correct response after $k$ sampling attempts, we supplement them with a golden rationale, ensuring that every query has at least one practically correct response and effectively rebalancing the long-tail distribution. We use hollow markers to represent the performance with tail data. As shown in Figure 2, across different numbers of sampling and model variants, the performance represented by the dashed line exceeds that of the solid line representing vanilla self-improve. The results demonstrate that rebalanced data helps mitigate performance degradation under a limited sampling budget. Moreover, a larger budget alleviates performance plateaus and boosts model efficacy to some extent. This study highlights the value of solutions to challenging queries in overcoming the limits of finite and progressively biased sampling.
@@ 5 Guided Self-Improvement
To mitigate the observed tail-narrowing phenomenon, a straightforward solution is to increase the sampling budget for tail data. However, directly tackling these difficult queries from scratch often results in low success rates and higher costs due to repeated failed attempts. Drawing from Socratic-style education (Chang, 2023; Dong et al., 2023b), we introduce guidance-based exploration techniques to improve sampling efficiency, such as learning from demonstrations (Schaal, 1996; Subramanian et al., 2016). We provide the model with tailored assistance in structured contexts, enabling it to address difficult queries more effectively. The following paragraphs outline four guiding strategies we propose.
@@ 6 Experiments
Table 1: Main results on six math reasoning tasks. The best result for each dataset is highlighted in bold, while the second-best result is marked with underline. Thm.QA denotes the TheoremQA task. Coverage refers to the number of unique problems solved in the Generate Step, while Sample Budget indicates the number of sampling during this step. The baselines include SFT, the vanilla Self-Improve ($k=8$), and brute-force sampling methods with higher computational demands: Self-Improve ($k=64$) and Self-Improve ($k=128$).
@@ 6.1 Experimental Setups
Baselines
To evaluate the impact of GSI on distributional adjustments, we compare it against SFT and Self-improve variants. We also include baselines with significantly higher sampling trails than GSI.
@@
  • Self-Improve ($k=64$ or $k=128$): Based on Vanilla Self-Improve ($k=8$), we add distribution re-balancing stage for tail data in each iteration. It performs brute-force sampling of tail-end data up to $k$ times without additional guidance.
@@ 6.2 Main Results
Compared to SFT, vanilla self-improve with $k=8$ boosts reasoning on held-out datasets but shows only marginal gains, with performance bottlenecks and degradation on held-in datasets. This aligns with our observation (§ 4), where we identify tail narrowing as the primary cause. To mitigate this, Self-Improve ($k=64$ or $k=128$) performs additional sampling on tail data, which re-balances the distribution. This adjustment significantly enhances both coverage and overall performance, with further improvements as the number of sampling increases. For example, in Llama2-7B, coverage increases from $52.6$% to $80.7$%, with performance improving from $21.53$ to $23.84$. In Mistral-7B, the rebalancing also reverses the observed performance decline. Thus, incorporating a resampling stage for challenging, heavy-tailed data proves essential.
@@
As the resampling increases from $64$ to $128$, the performance gains become marginal despite additional computational costs. This may be due to the model getting trapped in vast search space, especially on more challenging queries. In contrast, our strategy GSI, which leverages Socratic-style guidance, offers a more compute-efficient sampling. It outperforms brute-force sampling while utilizing only one-third of the sampling budget. Specifically, on Llama3-8B, the state reset strategy performs $0.32$M sampling, which is only one-third of the budget required for brute-force sampling. Moreover, the model shows improved held-in performance from $37.55$ to $41.64$ and generalizes well to held-out datasets.

### v2

@@ Abstract
Yiwen Ding1thanks:  Equal contribution. Work done during internship at Meituan. Corresponding author., Zhiheng Xi [^1] footnotemark: 1, Wei He1, Zhuoyuan Li5, Yitao Zhai2, Xiaowei Shi2, Xunliang Cai2, Tao Gui3†, Qi Zhang1,4, Xuanjing Huang1,4† 1 School of Computer Science, Fudan University   2 Meituan 3 Institute of Modern Languages and Linguistics, Fudan University 4 Key Laboratory of Intelligent Information Processing, Fudan University 5 Macau University of Science and Technology {ywding23,zhxi22}@m.fudan.edu.cn, {tgui, xjhuang}@fudan.edu.cn
@@ 1 Introduction
To address this imbalance, a common approach is to allocate more sampling trials to under-sampled, challenging queries (Tong et al., 2024), but this can be considerably more costly. In this paper, we propose Guided Self-Improvement (GSI), an efficient method that leverages interactive guidance signals inspired by the Socratic method (Chang, 2023; Dong et al., 2023b). Specifically, our approach introduces an additional resampling phase, termed distribution re-balancing for difficult queries, which is applied after the generation step in the self-improvement process. During this phase, we incorporate targeted guidance signals derived from oracle answers, stepwise rationales, and strong teacher supervision. These signals help to narrow the sampling space, reduce sampling difficulty, and minimize hallucinations during reasoning (Xie et al., 2023; Xi et al., 2024). As a result, GSI enables more effective exploration, increases solution coverage for challenging queries (Bansal et al., 2024), and mitigates the issue of tail narrowing during the sampling process.
@@ 2.1 Self-improvement for LLMs
Once filtered, the high-quality data supports post-training through methods like SFT (Zelikman et al., 2022; isheep) or preference-based techniques like Direct Preference Optimization (DPO, Yuan et al., 2024). This process is often iterative, allowing models to continually generate new data, filter it, and use it to refine their performance further (Zelikman et al., 2022; Gülçehre et al., 2023; Yuan et al., 2024).
@@ 3.1 Formulation of Self-improvement
Given a large language model $M_{0}$ and the original training dataset $\mathcal{D}=\{(x_{i},r_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is the problem, $r_{i}$ is the chain-of-thought rationale (Wei et al., 2022b) and $y_{i}$ represents the final answer. Each rationale $r_{i}$ consists of several intermediate steps, i.e., $r_{i}=[r_{i,1},\dots,r_{i,L}]$, where $L$ denotes the number of steps. The self-improvement process enhances the model’s reasoning ability through iterative refinement over $T$ cycles. Each iteration $t\in[1,T]$ consists of two main steps: Generate and Improve.
@@
At iteration $t$, the previous model $M_{t-1}$ generates multiple reasoning paths for each problem. Specifically, we allocate $k$ sampling times to each query $x_{i}\in\mathcal{D}$:
@@
The newly generated data points form a set $\mathcal{D}^{{}^{\prime}}=\{(x_{i},\hat{r}_{i}^{j},\hat{y}_{i}^{j})\mid x_{i}\in\mathcal{D},j=[1,k]\}$. Each candidate solution $\hat{y}_{i}^{j}$ is evaluated by a binary reward function $\textrm{rf}(y_{i},\hat{y_{i}})\in\{0,1\}$, which verifies correctness based on ground-truth answers $y_{i}$. Only correct solutions with $\textrm{rf}(\cdot)=1$ are filtered to form the high-quality dataset $\mathcal{D}_{t}$ for the current iteration $t$.
@@
In the $t$-th iteration, the model $M_{t}$ is fine-tuned on the self-generated high-quality dataset $\mathcal{D}_{t}$. The fine-tuning objective is to minimize the negative log-likelihood (NLL) loss:
@@ 4 Performance Bottleneck and Tail Narrowing in Self-Improvement
Despite extensive research into self-improvement methods (Gülçehre et al., 2023; Singh et al., 2024), the performance dynamics across successive training iterations remain underexplored. To this end, we conduct experiments across four backbone models to investigate the effects of sampling and tail data during iterative training. Tail data refers to samples for which the model rarely generates correct solutions during sampling. These samples lie in the “long tail" of data distributions.
@@
To uncover the relationship between performance trends with sampling, we vary the number of sampling $k$. Figure 2 reveals the following: (1) Lower sampling times (e.g., $k=1$ or $k=2$) degrade model performance, leading to negative gains. Llama2-7B, Mistral-7B, and Llama3-8B underperform compared to SFT on the original dataset (iteration 1). This decline stems from the models’ weaker reasoning abilities, which limit their coverage of challenging queries. Consequently, the models tackle only basic problems, resulting in a degradation or collapse of reasoning ability (Shumailov et al., 2023; Dohmatob et al., 2024a). (2) With larger sampling numbers $k$, performance improves across multiple iterations. The most notable gains occur during the first iteration, as shown by the green and orange solid lines in Figure 2. However, after the third iteration, progress halts, eventually reaching a performance plateau.
@@
As previously discussed, merely scaling the number of sampling eventually encounters performance bottlenecks. However, challenging low-probability samples, often referred to as tail data, are often considered more crucial for improving model performance (Sorscher et al., 2022; Liu et al., 2024; Tong et al., 2024).
To investigate the impact of tail data, we conduct additional experiments targeting these difficult examples. For queries that do not yield a correct response after $k$ sampling attempts, we supplement them with a golden rationale, ensuring that every query has at least one practically correct response and effectively rebalancing the long-tail distribution. We use hollow markers to represent the performance with tail data. As shown in Figure 2, across different numbers of sampling and model variants, the performance represented by the dashed line exceeds that of the solid line representing vanilla self-improve. The results demonstrate that rebalanced data helps mitigate performance degradation under a small number of sampling attempts (e.g., $k=1$ or $k=2$). Moreover, as $k$ increases, a larger number of sampling alleviates performance plateaus and boosts model efficacy to some extent. This study highlights the value of solutions to challenging queries in overcoming the limits of finite and progressively biased sampling.
@@ 5 Guided Self-Improvement
To mitigate the observed tail-narrowing phenomenon, a straightforward solution is to increase the sampling trials for tail data. However, directly tackling these difficult queries from scratch often results in low success rates and higher costs due to repeated failed attempts. Drawing from Socratic-style education (Chang, 2023; Dong et al., 2023b), we introduce guidance-based exploration techniques to improve sampling efficiency, such as learning from demonstrations (Schaal, 1996; Subramanian et al., 2016). We provide the model with tailored assistance in structured contexts, enabling it to address difficult queries more effectively. The following paragraphs outline four guiding strategies we propose.
@@ 6 Experiments
Table 1: Main results on six math reasoning tasks. The best result for each dataset is highlighted in bold, while the second-best result is marked with underline. Results marked in blue indicate average scores. Thm.QA denotes the TheoremQA task. Coverage refers to the number of unique problems solved in the Generate Step, while Sample Budget indicates the total number of sampling times during this step. “SI” refers to Self-Improve. The baselines include SFT, the vanilla Self-Improve, and Brute-Force Self-Improve.
@@ 6.1 Experimental Setups
Baselines.
To evaluate the impact of GSI on distributional adjustments, we compare it against SFT and Self-improve variants. We also include baselines with significantly higher sampling trials than GSI.
@@
  • Brute-Force Self-Improve ($k=64$ or $k=128$): Based on Vanilla Self-Improve ($k=8$), we add a distribution re-balancing stage for tail data in each iteration. It performs non-guided brute-force sampling of tail-end data up to $k$ times without additional guidance.
@@ 6.2 Main Results
Compared to SFT, vanilla self-improve with $k=8$ boosts reasoning on held-out datasets but shows only marginal gains, with performance bottlenecks and degradation on held-in datasets. This aligns with our observation (§ 4), where we identify tail narrowing as the primary cause. To mitigate this, Brute-Force Self-Improve ($k=64$ or $k=128$) performs additional sampling on tail data, which re-balances the distribution. This adjustment significantly enhances both coverage and overall performance, with further improvements as the number of samples increases. For example, in Llama2-7B, coverage increases from $52.6$% to $80.7$%, with performance improving from $21.53$ to $23.84$. In Mistral-7B, the rebalancing also reverses the observed performance decline. Thus, incorporating a resampling stage for challenging, heavy-tailed data proves essential.
@@
To optimize the amount of sampling computation required for self-improvement, we analyze the sampling efficiency during the Generate step phase. For each query $x_{i}\in\mathcal{D}$ at iteration $t$, we perform $k_{i,t}$ sampling operations. The cumulative sample budget across all queries and $T$ iterations is defined as: Sample $\text{Sample Budget}=\sum_{x_{i}\in\mathcal{D}}\sum_{t=1}^{T}k_{i,t}$.
When scaling resampling operations from 64 to 128, we observe that the improvement in sampling coverage becomes slower, and the performance gains diminish despite the additional computational costs. This suggests that the model may get trapped in vast search space, particularly when dealing with more challenging queries. In contrast, our strategy GSI, which leverages Socratic-style guidance, achieves more compute-efficient sampling. It outperforms brute-force sampling while using only one-third of the sampling budget. Specifically, on Llama3-8B, the state reset strategy performs $0.32$M sampling, which is only one-third of the budget required for brute-force sampling. Moreover, the model shows improved held-in performance from $37.55$ to $41.64$ and generalizes well to held-out datasets.
@@ 7.3 Short-cutting in Generated Rationales
7.4 Impact of the Sampling Hyperparameter
We investigate the impact of the hyperparameter $k$, which determines the number of sampling times allocated to each query. As shown in Table 3, increasing $k$ yields substantial performance improvements. Notably, the most significant improvements occur up to $k=8$, after which the gains begin to plateau. Therefore, we choose $k=8$ as a practical configuration, achieving a balanced trade-off between computational cost and performance gains.
@@ 8 Conclusion
Models Setting AQuA GSM8K MATH
Llama3-8B $k=2$ $40.95$ $56.94$ $13.20$
$k=4$ $42.91$ $61.49$ $15.28$
$k=8$ $46.46$ 62.62 15.84
$k=16$ 47.24 $62.32$ $15.64$
Mistral-7B $k=2$ $29.92$ $40.10$ $6.82$
$k=4$ $31.49$ $47.46$ $8.76$
$k=8$ $33.07$ 50.95 $9.68$
$k=16$ 33.46 $50.19$ 9.82
Table 3: Performance of GSI (State Reset) with varying values of $k$.
@@
Acknowledgment
The authors wish to thank the anonymous reviewers for their helpful comments. This work was supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100), National Natural Science Foundation of China (No. 62476061,62206057,62076069), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500), Program of Shanghai Academic Research Leader under grant 22XD1401100. This research was also supported by Meituan.